Please help! How do I solve this rational equation?
2/x+5 - 3/x+3=1/x

Answer by kapilsinghi(68)

2/x + 5 - 3/x + 3 =1/x
-1/x + 8 = 1/x
8 = 2/x
x = 1/4

kapilsinghi123@gmail.com